OpenStudy (anonymous):

use distributive property to simply this expression: (6+9v)6 please explain im so confused!!

jimthompson5910 (jim_thompson5910):

(6+9v)6 6(6+9v) 6*6 + 6*9v I'll let you finish up

jimthompson5910 (jim_thompson5910):

In general (b+c)a = a(b+c) = ab+ac

OpenStudy (anonymous):

ty so much!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

This really confusing